# Effect of spin in binary neutron star mergers

### v1

@@ 3.7.3 Spin
We have $7$ configurations that form BHs, $6$ of which experience prompt collapse and the other one experiences a delayed collapse. The apparent horizons of the black holes are detected with AHFinderDirect, see Thornburg (1996, 2004); Brown et al. (2009). The quasi-local mass and the dimensionless spin parameter of the BH at time $t_{dyn,1}$ for all binaries experience prompt collapse and at time $t_{dyn,2}$ for the case that experiences delay collapse to the BH, are presented in Table LABEL:bh_properties. We also list the time difference between the BH formation and the merger, $t_{BH}-t_{merger}$. The BH remnant formed from binary neutron star mergers of the aligned spinning case with mass $M_{tot}=4.10M_{\odot}$ is the fastest spinning BH with $\chi_{BH}=0.92$, whereas for the irrotational case the BH spin value is $\sim 0.83$.
Table 4: The quasi-local mass and spin of the BH remnant at formation time $t_{form}$ and at final times. The time difference $t_{BH}-t_{merger}$ shows the BH formation time, where negative or positive sign indicate that the BH formed before or after the merger, respectively.
@@ 3.8 Disc structure
In this section, we discuss the distribution, composition and velocity of the matter in the disc, at time $t_{dyn,1}$, for the different total mass cases separately.
@@
The comparison of the two aligned spinning cases $\chi=0.4$ and $0.67$, shows the presence of two spiral arms, see Fig.2. The extent of the arm increases with amount of spin. The inner parts of the arms cover the remnant and mass is transferred from the arms to the region with rest mass density $\rho=10^{11}gcm^{-3}$ (see Fig.8). The velocity of the matter reaches $\sim 0.20$c at the place where the feeding stream hits the region with density $\rho=10^{11}$$gcm^{-3}$. The composition of the spiral arms is very neutron rich with $Y_{e}\leq 0.05$, see Figs. 9 and 3. In the inner parts of the disc the electron fraction increases to $0.07$ and $0.12$, at most, for the two cases $\chi=0.4$ and $0.67$, respectively.
@@ 4 Summary and Conclusion
We study the effect of the spin in BNS mergers. We consider $17$ configurations with $3$ different total masses and having both equal and unequal mass. Also spin configurations are varied, to cover cases where both stars’ spins are aligned, or anti-aligned or mixed, with respect to the orbital spin. We investigate the impact of the initial spin on gravitational radiation emission, on properties of ejected matter and thermodynamic properties of all these systems as well as gravitational wave strains, their detectability, the maximum temperature and the rest-mass density of matter, the neutrino energy and their luminosities.
@@
The presence of spin changes the structure of the NSs during the merger. The formation of tidal tails or spiral arms is not seen in the equal mass irrotational BNS mergers. Four of our configurations for $M_{tot}=2.55M_{\odot}$ have these features due to spin. We see the formation of spiral arms in models $\chi=0.4$ and $0.67$ and of tidal tails in models $\chi=0.2$, $\chi=0[2]$, and $0[3]$. The extent of the tidal tail for $\chi=0.2$ is very small and the tail quickly dissolves. However, the spiral arms whose strength increases as the spin increases, extend to very large distances, for models $\chi=0.4$ and $0.67$ cases. The composition of the spiral arms and tidal tails are very neutron rich. This is a key parameter to determine the heavy r-process nucleosynthesis Foucart (2020).
@@
The values of spin investigated in this study are higher than what is observed and we have shown the impact of such high spins on gravitational waves emission and properties of ejected matter. Hence, it is possible to distinguish them by gravitational wave observations. The differences in possible electromagnetic counterparts was not investigated here.
@@ Acknowledgements
RM and BK would like to thank Ian Hawke for his comments/suggestions, guidance and support. BK would like to thank Roland Haas for his constant support, guidance and his useful comments/suggestions on every part of the study, David Radice for his useful comments and guidance, Wolfgang Tichy for his support and guidance, and Zachariah B. Etienne for his valuable comments.
@@
This work used TACC Stampede at XSEDE through allocation PHY160053 from the Extreme Science and Engineering Discovery Environment (XSEDE), which was supported by National Science Foundation grant number #1548562. The work has been performed under the Project HPC-EUROPA3 (INFRAIA-2016-1-730897), with the support of the EC Research Innovation Action under the H2020 Programme; in particular, BK gratefully acknowledges the support of The Edinburgh University and the computer resources and technical support provided by EPCC. The numerical calculations reported in this paper were partially performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).

### v2

@@ 3.7.3 Spin
We have $7$ configurations that form BHs, $6$ of which experience prompt collapse and the other one experiences a delayed collapse. The apparent horizons of the black holes are detected with AHFinderDirect, see Thornburg (1996, 2004); Brown et al. (2009). The quasi-local mass and the dimensionless spin parameter of the BH at time $t_{dyn,1}$ for all binaries experience prompt collapse and at time $t_{dyn,2}$ for the case that experiences delay collapse to the BH, are presented in Table LABEL:bh_properties. We also list the time difference between the BH formation and the merger, $t_{BH}-t_{merger}$. The delay in BH formation for prompt collapse models due to spin is also reported in a recent study by  Schianchi et al. (2024). The BH remnant formed from binary neutron star mergers of the aligned spinning case with mass $M_{tot}=4.10M_{\odot}$ is the fastest spinning BH with $\chi_{BH}=0.92$, whereas for the irrotational case the BH spin value is $\sim 0.83$.
Table 4: The quasi-local mass and spin of the BH remnant at formation time $t_{form}$ and at final times $t_{f}$. The time difference $t_{BH}-t_{merger}$ shows the BH formation time, where negative or positive sign indicate that the BH formed before or after the merger, respectively.
@@ 3.8 Disc structure
In this section, we discuss the distribution, composition and velocity of the matter in the disc, at time $t_{dyn,1}$, for the different total mass cases separately. We also present the impact of initial spin on disc masses.
@@
The comparison of the two aligned spinning cases $\chi=0.4$ and $0.67$, shows the presence of two spiral arms, see Fig.2. The extent of the arm increases with amount of spin. The inner parts of the arms cover the remnant and mass is transferred from the arms to the region with rest mass density $\rho=10^{11}gcm^{-3}$ (see Fig.8). The velocity of the matter reaches $\sim 0.20$c at the place where the feeding stream hits the region with density $\rho=10^{11}$$gcm^{-3}$. The composition of the spiral arms is very neutron rich with $Y_{e}\leq 0.05$, see Figs. 3 and 9. In the inner parts of the disc the electron fraction increases to $0.07$ and $0.12$, at most, for the two cases $\chi=0.4$ and $0.67$, respectively.
@@
3.8.4 Impact of spin on disc mass
In this section, we discuss the effect of initial effective spin on the disc mass. In order to calculate the baryonic mass of the disc, we integrate the mass with a rest-mass density below $\rho<10^{13}$$gcm^{-3}$, similarly to  Radice et al. (2018) and references therein, see also Camilletti et al. (2024), within a radius of $\sim 443$km. We calculate ejected matter properties therein. If a BH is formed, we remove the region with lapse function values lower than $0.3$. The disc masses are given in Table LABEL:ejecta_properties_table, and the dependence of disc mass on the initial effective spin is shown in Fig. 10. In this figure, we provide the values of disc mass at time $t_{dyn,2}$, if the mass exists at that point in time. In Fig.11 we show linear fits of the correlation between the sum of the individial spin magnitudes and disc mass, at time $t_{dyn,1}$ Although a correlation is clearly present, the fits display a big deviation due to low resolution. The fits are produced using linear regression through the least squares method.
Figure 10: Effect of initial effective spin, $\chi$, on the disk mass at $t_{dyn,2}$, if the disc mass exists at that time. Binaries with the total mass of $M_{tot}=2.55M_{\odot}$, $3.05M_{\odot}$, and $4.10M_{\odot}$ are represented by black squares, red diamonds and green circles, respectively. The equal mass, irrotational cases are shown by triangles with colour corresponding to the total mass. The disc masses plotted for $M_{tot}=2.55M_{\odot}$ have a resolution of $h=307$m. The most massive disc is formed for the case $\chi=0.4$ with a mass of $\sim 0.36M_{\odot}$.
For models with $M_{tot}=2.55M_{\odot}$, we provide the disc masses at $t_{dyn,1}$ for $\chi=0[1]$ and $\chi=0.2$, and at $t_{dyn,2}$ for other models. For the latter models, disc masses decrease between times $t_{dyn,1}$ to $t_{dyn,2}$ for $\chi=-0.65$, $\chi=0[3]$, and $\chi=0.67$ cases, but increase for all other models. At $t_{dyn,1}$, all models except for $\chi=-0.4$ produce higher disc mass as compared to the irrotational model. In addition to that, aligned spinning models result in higher disc mass as compared to the anti-aligned spinning models with the same spin magnitude. However, both anti-aligned (aligned) spinning models produce a more (less) massive disc as the spin magnitude increases. We also see that, even though the effective spin is zero for three cases, $\chi=0[1]$, $0[2]$ and $0[3]$, individual spins are important and the disc mass increases as the individual spin magnitude increases. The most massive disc is formed for the $\chi=0.4$ model and reaches $\sim 0.36M_{\odot}$. The dependence of initial spin on disc mass, assuming one component is spinning, has also been presented by Papenfort et al. (2022); Rosswog et al. (2023, 2024).
For models with $M_{tot}=3.05M_{\odot}$, the comparison of equal and unequal mass irrotational models shows that even though they result in prompt black hole formation, the disc for the unequal mass model with $q=2.05$ is much more massive. It has a mass of $\sim 0.18M_{\odot}$. Between times $t_{dyn,1}$ to $t_{dyn,2}$ disc masses decrease by $\sim 0.02M_{\odot}$ and $\sim 0.03M_{\odot}$, for equal and unequal mass spinning models, respectively. The most massive disc formed at dyn,2 is observed for the equal mass spinning model with a mass of $\sim 0.27M_{\odot}$. This is despite being simulated for a shorter time as compared to the unequal mass, spinning model.
For models with $M_{tot}=4.10M_{\odot}$ the merger of the anti-aligned spinning model results in a disc with negligible mass, whereas both-aligned spinning model results in forming a much more massive disc as compared to the irrotational case. Its mass is $\sim 0.36M_{\odot}$.
Figure 11: Linear fits of correlation between the sum of individial spin magnitude $\left|\chi_{1}\right|$+$\left|\chi_{2}\right|$ and disc mass, at $t_{dyn,1}$.
@@ 4 Summary and Conclusion
We study the effect of the spin in BNS mergers. We consider $17$ configurations with $3$ different total masses and having both equal and unequal mass. Also spin configurations are varied, to cover cases where both stars’ spins are aligned, or anti-aligned or mixed, with respect to the orbital spin. We investigate the impact of the initial spin on gravitational radiation emission, on properties of ejected matter and thermodynamic properties of all these systems as well as gravitational wave strains, their detectability, the maximum temperature and the rest-mass density of matter, the neutrino energy and their luminosities, and disc masses.
@@
The presence of spin changes the structure of the NSs during the merger. The formation of tidal tails or spiral arms is not seen in the equal mass irrotational BNS mergers. Four of our configurations for $M_{tot}=2.55M_{\odot}$ have these features due to spin. We see the formation of spiral arms in models $\chi=0.4$ and $0.67$ and of tidal tails in models $\chi=0.2$, $\chi=0[2]$, and $0[3]$. The extent of the tidal tail for $\chi=0.2$ is very small and the tail quickly dissolves. However, the spiral arms whose strength increases as the spin increases, extend to very large distances, for models $\chi=0.4$ and $0.67$ cases. The composition of the spiral arms and tidal tails are very neutron rich. This is a key parameter to determine the heavy r-process nucleosynthesis Foucart (2020). Our results of the effect of spin on the tidal disruption of a neutron star in equal mass binary neutron star mergers is in agreement with the studies of  Papenfort et al. (2022); Rosswog et al. (2023, 2024) where only one component has a spin.
@@
We present the dependence of disc mass on the initial spin. Our simulations produce discs with mass $\sim 0.36M_{\odot}$ in cases of low (high) mass binary neutron star mergers, despite the latter promptly collapses to a BH. Even though both equal and unequal mass irrotational cases with $M_{tot}=3.05M_{\odot}$ promptly collapse to a BH, the latter form a massive disc with mass $\sim 0.18M_{\odot}$. This elucidates the dependence of mass ratio on the disc mass. As the equal mass, spinning case with $M_{tot}=3.05M_{\odot}$ experiences a hypermassive neutron star phase before finally collapsing to a BH, it has the most massive disc with mass $\sim 0.27M_{\odot}$. It is noteworthy to point out that mergers of equal high mass and rapidly spinning neutron stars yield a substantially massive disc with a mass of $\sim 0.36M_{\odot}$. This shows the importance of considering high spins on electromagnetic counterparts even for equal mass binaries that promptly collapse to a BH. Since disc mass is the quantity with the most impact on the kilonova peak luminosity, the correlation in Fig.11 suggests that one could use EM data to constrain $\left|\chi_{1}\right|$+$\left|\chi_{2}\right|$, breaking the degeneracy in the measurement of spin from gravitational wave alone.  [^1] David Radice, personal communication
The values of spin investigated in this study are higher than what is observed and we have shown the impact of such high spins on gravitational waves emission and properties of ejected matter. Hence, it is possible to distinguish them by gravitational wave observations. The differences in possible electromagnetic counterparts are not investigated here.
@@ Acknowledgements
RM and BK would like to thank Ian Hawke for his comments/suggestions, guidance and support. BK would like to thank Roland Haas for his constant support, guidance and his useful comments/suggestions on every part of the study, David Radice for his many useful comments/suggestions, help and guidance, Wolfgang Tichy for his support and guidance, and Zachariah B. Etienne for his valuable comments.
@@
This work used TACC Stampede at XSEDE through allocation PHY160053 from the Extreme Science and Engineering Discovery Environment (XSEDE), which was supported by National Science Foundation grant number #1548562. The work has been performed under the Project HPC-EUROPA3 (INFRAIA-2016-1-730897), with the support of the EC Research Innovation Action under the H2020 Programme; in particular, BK gratefully acknowledges the support of The University of Edinburgh and the computer resources and technical support provided by EPCC. The numerical calculations reported in this paper were partially performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).